# Multi-intention Inverse Q-learning for Interpretable Behavior Representation

### v3

@@ 6 Conclusion
This paper proposes the class of EM-based hierarchical inverse Q-learning (HIQL) algorithms, extending the IQL algorithm [KHWB20] for multi-intention IRL problems. Empirical results demonstrate that HIQL outperforms the state-of-the-art on both empirical and real-world datasets, and is able to produce interpretable behavior characteristics. We also mathematically characterized typical exploration behavior of rodents during value-based decision-making using HIQL. As future work, we plan to relax the Markov assumption about the intention transition dynamics in HIQL to incorporate non-Markovian intention dynamics, and also extend HIQL for an unknown number of reward functions [DR12, MH12, SS14].
@@ Acknowledgments
This work has been funded as part of BrainLinks-BrainTools, which is funded by the Federal Ministry of Economics, Science and Arts of Baden-Württemberg within the sustainability programme for projects of the Excellence Initiative II; as well as the Bernstein Award 2012, the Research Unit 5159 “Resolving Prefrontal Flexibility” (Grant DI 1908/11-1), and the Deutsche Forschungsgemeinschaft (Grants DI 1908/3-1, DI 1908/11-1, and DI 1908/6-1), all to I.D.
Appendix A Proof of Theorem 4.3
@@
$J\left(\Theta^{+}\mid\Theta\right)=\mathbf{E}_{\xi\sim\mathcal{D},\eta}\left[\log\mathbf{P}\left(\xi,\eta\mid\Theta^{+}\right)\right]$
$\quad=\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)\log\mathbf{P}\left(\xi,\eta\mid,\Theta^{+}\right)\right]$
$\quad=\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)\log\left(\Pi^{+}_{z_{0}}\pi_{r_{z_{0}}^{+}}(s_{0},a_{0})\prod_{t=1}^{n}\Lambda^{+}_{z_{t-1}z_{t}}\pi_{r_{z_{t}}^{+}}(s_{t},a_{t})\right)\right]$
$\quad=\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)\log\Pi^{+}_{z_{0}}\right]+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)\sum_{t=1}^{n}\log\Lambda^{+}_{z_{t-1}z_{t}}\right]$
$\qquad+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)\sum_{t=0}^{n}\log\pi_{r_{z_{t}}^{+}}(s_{t},a_{t})\right]$
$\quad=\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\underbrace{\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)I_{i}(z_{0})}_{\mathbf{P}(z_{0}=i\mid\xi,\Theta)}\log\Pi^{+}_{i}\right]+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\underbrace{\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)I_{i}(z_{t-1})I_{j}(z_{t})}_{\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)}\log\Lambda^{+}_{ij}\right]$
$\qquad+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{t=0}^{n}\underbrace{\sum_{\eta}\mathbf{P}(\eta\mid\xi,\Theta)I_{i}(z_{t})}_{\mathbf{P}(z_{t}=i\mid\xi,\Theta)}\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]$
$\quad=\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathbf{P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]$ (A.1a)
$\qquad+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]$ (A.1b)
$\qquad+\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{t=0}^{n}\mathbf{P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right],$ (A.1c)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Pi^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathbf{P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]\\ \mbox{\rm subject to}&\Pi^{+}\succeq 0,\ \mathbf{1}^{T}\Pi^{+}=1,\end{array}$ (3)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Lambda^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]\\ \mbox{\rm subject to}&\Lambda^{+}_{i:}\succeq 0,\ \mathbf{1}^{T}\Lambda^{+}_{i:}=1,\quad i=1,\ldots,K.\end{array}$ (4)
Note that (A.1c) has similar structure as the optimization objective of IRL (problem (1)), and can be further separated for each $r\in\mathcal{R}$, thus maximizing (A.1c) over $\mathcal{R}^{+}$ is equivalent to independently solving an IRL problem for each $r_{i}$, with the $t$th demonstration weighted by $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$:
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $r_{i}^{+}$)}&\mathbf{E}_{\xi\sim\mathcal{D}}\left[\sum_{t=0}^{n}\mathbf{P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]\\ \mbox{\rm subject to}&\mbox{\rm$\pi_{r_{i}^{+}}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{\rm$Q(s,a)=r_{i}^{+}(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}.\end{array}$ (5)
@@
Appendix B Computing required posterior probabilities
To obtain the posterior probability terms $\mathbf{P}(z_{0}=i\mid\xi,\Theta)$, $\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$, and $\mathbf{P}(z_{t}=i\mid\xi,\Theta)$ required for evaluating the objective functions of (3), (4) and (5), we apply the Baum-Welch algorithm [BP66] as follows.
@@ Appendix B Computing required posterior probabilities
$\alpha_{t}(i)$ $=\mathbf{P}(\xi_{0:t},z_{t}=i\mid\Theta)$ (B.1)
=$=\left\{\begin{array}[]{ll}\Pi_{i}\pi_{r_{i}}(s_{0},a_{0})&t=0\\ \alpha_{t-1}^{T}P_{:i}\pi_{r_{i}}(s_{t},a_{t})&t>0,\end{array}\right.$ (B.4)
@@
$\beta_{t}(i)$ $=\mathbf{P}(\xi_{t+1:n}\mid z_{t}=i\mid\Theta)$ (B.5)
=$=\left\{\begin{array}[]{ll}\sum_{j=1}^{K}\beta_{t+1}(j)P(ij)\pi_{r_{j}}(s_{t+1},a_{t+1})&t<n\\ 1&t=n,\end{array}\right.$ (B.8)
for all $i=1,\ldots,K$. With (B.4) and (B.8), we can efficiently obtain the forward and backward probabilities for each demonstration in a recursive form. Finally, we can compute the posterior probabilities
$\mathbf{P}(z_{t}=i\mid\xi,\Theta)=\frac{\alpha_{t}(i)\beta_{t}(i)}{\alpha_{t}^{T}\beta_{t}},$ (B.9)
@@
$\mathbf{P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)=\frac{\alpha_{t}(i)P_{ij}\pi_{r_{j}}(s_{t+1},a_{t+1})\beta_{t+1}(j)}{\sum_{i=1}^{K}\sum_{j=1}^{K}\alpha_{t}(i)P_{ij}\pi_{r_{j}}(s_{t+1},a_{t+1})\beta_{t+1}(j)}.$ (B.10)

### v4

@@ 6 Conclusion
This paper proposes the class of EM-based hierarchical inverse Q-learning (HIQL) algorithms, extending the IQL algorithm (Kalweit et al., 2020) for multi-intention IRL problems. Empirical results demonstrate that HIQL outperforms the state-of-the-art on both synthesized and real-world datasets, and is able to produce interpretable behavior characteristics. We also mathematically characterized typical exploration behavior of rodents during value-based decision-making using HIQL.
Compared to the state-of-the-art algorithm for characterizing animal behavior, DIRL (Ashwood et al., 2022a), the advantages of our HIQL algorithm are two fold. First, the assumptions about the underlying intention transition dynamics in HIQL align better with those observed in real-world behavioral experiments that animal and human alternate between discrete strategies during decision making (Ashwood et al., 2022b), rather than optimize their policy over a continuous time-varying reward as assumed by DIRL. Second, the inner-loop IRL problem solver adapted by HIQL, the class of IQL algorithms, were shown to have better performance in learning the expert’s unknown reward and higher computing efficiency than the maximum entropy IRL algorithm used by DIRL (Kalweit et al., 2020).
As future work, we plan to relax the Markov assumption about the intention transition dynamics in HIQL to incorporate non-Markovian intention dynamics, and also extend HIQL for an unknown number of reward functions (Dimitrakakis & Rothkopf, 2012; Michini & How, 2012; Surana & Srivastava, 2014).
@@ Acknowledgments
This work has been funded as part of BrainLinks-BrainTools, which is funded by the Federal Ministry of Economics, Science and Arts of Baden-Württemberg within the sustainability programme for projects of the Excellence Initiative II; as well as the Bernstein Award 2012, the Research Unit 5159 “Resolving Prefrontal Flexibility” (Grant DI 1908/11-1), and the Deutsche Forschungsgemeinschaft (Grants DI 1908/3-1 and DI 1908/6-1) to I.D., and CRC/TRR 384 “IN-Code” to I.D. and J.B.
Appendix A Theoretical and technical details
A.1 Proof of Theorem 4.3
@@ Appendix A Theoretical and technical details
$J\left(\Theta^{+}\mid\Theta\right)=\mathop{\bf E}_{\xi\sim\mathcal{D},\eta}\left[\log\mathop{\bf P}\left(\xi,\eta\mid\Theta^{+}\right)\right]$
$\quad=\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)\log\mathop{\bf P}\left(\xi,\eta\mid,\Theta^{+}\right)\right]$
$\quad=\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)\log\left(\Pi^{+}_{z_{0}}\pi_{r_{z_{0}}^{+}}(s_{0},a_{0})\prod_{t=1}^{n}\Lambda^{+}_{z_{t-1}z_{t}}\pi_{r_{z_{t}}^{+}}(s_{t},a_{t})\right)\right]$
$\quad=\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)\log\Pi^{+}_{z_{0}}\right]+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)\sum_{t=1}^{n}\log\Lambda^{+}_{z_{t-1}z_{t}}\right]$
$\qquad+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)\sum_{t=0}^{n}\log\pi_{r_{z_{t}}^{+}}(s_{t},a_{t})\right]$
$\quad=\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\underbrace{\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)I_{i}(z_{0})}_{\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)}\log\Pi^{+}_{i}\right]+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\underbrace{\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)I_{i}(z_{t-1})I_{j}(z_{t})}_{\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)}\log\Lambda^{+}_{ij}\right]$
$\qquad+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{t=0}^{n}\underbrace{\sum_{\eta}\mathop{\bf P}(\eta\mid\xi,\Theta)I_{i}(z_{t})}_{\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)}\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]$
$\quad=\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]$ (A.1a)
$\qquad+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]$ (A.1b)
$\qquad+\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{t=0}^{n}\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right],$ (A.1c)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Pi^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)\log\Pi^{+}_{i}\right]\\ \mbox{\rm subject to}&\Pi^{+}\succeq 0,\ \mathbf{1}^{T}\Pi^{+}=1,\end{array}$ (7)
@@
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $\Lambda^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{i=1}^{K}\sum_{j=1}^{K}\sum_{t=1}^{n}\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)\log\Lambda^{+}_{ij}\right]\\ \mbox{\rm subject to}&\Lambda^{+}_{i:}\succeq 0,\ \mathbf{1}^{T}\Lambda^{+}_{i:}=1,\quad i=1,\ldots,K.\end{array}$ (8)
Note that (A.1c) has similar structure as the optimization objective of IRL (problem (1)), and can be further separated for each $r\in\mathcal{R}$, thus maximizing (A.1c) over $\mathcal{R}^{+}$ is equivalent to independently solving an IRL problem for each $r_{i}$, with the $t$th demonstration weighted by $\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$:
maximize (over $\begin{array}[]{ll}\mbox{\rm maximize (over $r_{i}^{+}$)}&\mathop{\bf E}_{\xi\sim\mathcal{D}}\left[\sum_{t=0}^{n}\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)\log\pi_{r_{i}^{+}}(s_{t},a_{t})\right]\\ \mbox{\rm subject to}&\mbox{\rm$\pi_{r_{i}^{+}}(s,a)=\exp\left(Q(s,a)-\log\sum\exp Q(s,\cdot)\right)$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}\\ &\mbox{\rm$Q(s,a)=r_{i}^{+}(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, for all $s\in\mathcal{S}$, $a\in\mathcal{A}$}.\end{array}$ (9)
@@
A.2 Computing required posterior probabilities
To obtain the posterior probability terms $\mathop{\bf P}(z_{0}=i\mid\xi,\Theta)$, $\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$, and $\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$ required for evaluating the objective functions of (7), (8) and (9), we apply the Baum-Welch algorithm (Baum & Petrie, 1966) as follows.
@@ A.2 Computing required posterior probabilities
$\alpha_{t}(i)$ $=\mathop{\bf P}(\xi_{0:t},z_{t}=i\mid\Theta)$ (A.2)
=$=\left\{\begin{array}[]{ll}\Pi_{i}\pi_{r_{i}}(s_{0},a_{0})&t=0\\ \alpha_{t-1}^{T}\Lambda_{:i}\pi_{r_{i}}(s_{t},a_{t})&t>0,\end{array}\right.$ (A.5)
@@
$\beta_{t}(i)$ $=\mathop{\bf P}(\xi_{t+1:n}\mid z_{t}=i,\Theta)$ (A.6)
=$=\left\{\begin{array}[]{ll}\sum_{j=1}^{K}\beta_{t+1}(j)\Lambda_{ij}\pi_{r_{j}}(s_{t+1},a_{t+1})&t<n\\ 1&t=n,\end{array}\right.$ (A.9)
for all $i=1,\ldots,K$. With (A.5) and (A.9), we can efficiently obtain the forward and backward probabilities for each demonstration in a recursive form. Finally, we can compute the posterior probabilities
$\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)=\frac{\alpha_{t}(i)\beta_{t}(i)}{\alpha_{t}^{T}\beta_{t}},$ (A.10)
@@
$\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)=\frac{\alpha_{t}(i)\Lambda_{ij}\pi_{r_{j}}(s_{t+1},a_{t+1})\beta_{t+1}(j)}{\sum_{i^{\prime}=1}^{K}\sum_{j^{\prime}=1}^{K}\alpha_{t}(i^{\prime})\Lambda_{i^{\prime}j^{\prime}}\pi_{r_{j^{\prime}}}(s_{t+1},a_{t+1})\beta_{t+1}(j^{\prime})}.$ (A.11)
Appendix B Algorithms
Algorithm 1 Inverse action-value iteration.
given expert demonstrations $\mathcal{D}$.
1:initialize $r$, $Q$.
2:Obtain expert policy $\pi^{\mathrm{E}}$ by counting the state-action visitation in $\mathcal{D}$.
3:repeat
4:    for all $s\in\mathcal{S}$ do
5:        Calculate $\mu(s,a)\coloneqq\log(\pi^{\mathrm{E}}(s,a))-\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$ for all $a\in\mathcal{A}$.
6:        Update $r(s,\cdot)$ by solving (5) via least squares.
7:        Update $Q(s,a)\coloneqq r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$ for all $a\in\mathcal{A}$.
8:    end for
9:until stopping criterion is satisfied.
Algorithm 2 Inverse Q-learning.
given expert demonstrations $\mathcal{D}$; learning rate $\alpha_{r}$, $\alpha_{Q}$, and $\alpha_{\mathrm{Sh}}$.
1:initialize $r$, $Q$, $Q^{\mathrm{Sh}}$, and state-action visitation counter $\rho$.
2:repeat
3:    for all $\xi\in\mathcal{D}$ do
4:        for $t=0,\ldots,n-1$ do
5:           Observe state-action pairs $(s_{t},a_{t},s_{t+1})$ from $\xi$, and update $\rho(s_{t},a_{t})\coloneqq\rho(s_{t},a_{t})+1$.
6:           Obtain $\pi^{\mathrm{E}}(s_{t},a)\coloneqq\rho(s_{t},a)/\sum_{a^{\prime}\in\mathcal{A}}\rho(s_{t},a^{\prime})$ for all $a\in\mathcal{A}$.
7:           Update $Q^{\mathrm{Sh}}(s_{t},a_{t})\coloneqq(1-\alpha_{\mathrm{Sh}})Q^{\mathrm{Sh}}(s_{t},a_{t})+\alpha_{\mathrm{Sh}}\cdot\gamma\max_{a\in\mathcal{A}}Q(s_{t+1},a)$.
8:           Calculate $\mu(s_{t},a)\coloneqq\log(\pi^{\mathrm{E}}(s_{t},a))-Q^{\mathrm{Sh}}(s_{t},a)$ for all $a\in\mathcal{A}$.
9:           Update $r(s_{t},a_{t})\coloneqq(1-\alpha_{r})r(s_{t},a_{t})+\alpha_{r}\left(\mu(s_{t},a_{t})+\frac{1}{\mathop{\bf card}(\mathcal{A}_{-a_{t}})}\sum_{b\in\mathcal{A}_{-a_{t}}}(r(s_{t},b)-\mu(s_{t},b))\right)$.
10:           Update $Q(s_{t},a_{t})\coloneqq(1-\alpha_{Q})Q(s_{t},a_{t})+\alpha_{Q}(r(s_{t},a_{t})+\gamma\max_{a\in\mathcal{A}}Q(s_{t+1},a))$.
11:        end for
12:    end for
13:until stopping criterion is satisfied.
Algorithm 3 Hierarchical inverse Q-learning.
given expert demonstrations $\mathcal{D}$ and reward set cardinality $K$.
1:initialize $\Pi$, $\Lambda$, $r_{1},\ldots,r_{K}$.
2:repeat
3:    Calculating the Boltzmann policy $\pi_{r_{i}}$ for all $i=1,\ldots,K$.
4:    for all $\xi\in\mathcal{D}$ do
5:        for $i=1,\ldots,K$, $j=1,\ldots,K$, $t=0,\ldots,n$ do
6:           Calculate $\alpha_{t}(i)$ and $\beta_{t}(i)$ according to (A.5) and (A.9).
7:           Calculate $\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$ and $\mathop{\bf P}(z_{t-1}=i,z_{t}=j\mid\xi,\Theta)$ according to (A.10) and (A.11).
8:        end for
9:    end for
10:    Update $\Pi$ and $\Lambda$ according to (10) and (11).
11:    for $i=1,\ldots,K$ do
12:        initialize expert demonstration subset $\mathcal{D}^{\prime}$.
13:        for all $\xi\in\mathcal{D}$ do
14:           for $t=0,\ldots,n-1$ do
15:               Observe state-action pairs $(s_{t},a_{t},s_{t+1})$ from $\xi$; sample $u\sim\mathcal{U}(0,1)$.
16:               if $u<\mathop{\bf P}(z_{t}=i\mid\xi,\Theta)$ then
17:                  $\mathcal{D}^{\prime}\coloneqq\mathcal{D}^{\prime}\cup\{(s_{t},a_{t},s_{t+1})\}$.
18:               end if
19:           end for
20:        end for
21:        Update $r_{i}$ by applying Algorithm 1 to $\mathcal{D}^{\prime}$ if $P$ is known, else using Algorithm 2.
22:    end for
23:until stopping criterion is satisfied.